# Scalable analysis of stop-and-go waves

### v1

@@ Abstract
Analyzing stop-and-go waves at the scale of miles and hours of data is an emerging challenge in traffic research. In the past, datasets were of limited scale and could be easily analyzed by hand or with rudimentary methods to identify a very limited set of traffic waves present within the data. The past 5 years have seen an explosion in the availability of large-scale traffic data containing traffic waves and complex congestion patterns, making existing approaches unsuitable for repeatable and scalable analysis of traffic waves in this data. This paper makes a first step towards addressing this shortcoming by introducing an automatic and scalable stop-and-go wave identification method capable of capturing wave generation, propagation, dissipation, as well as bifurcation and merging, which have previously been observed only very rarely. Using a concise and simple critical-speed based definition of a stop-and-go wave, the proposed method identifies all wave boundaries that encompass spatio-temporal points where vehicle speed is below a chosen critical speed. The method is built upon a graph-based representation of the spatio-temporal points associated with stop-and-go waves, specifically wave front (start) points and wave tail (end) points, and approaches the solution as a graph component identification problem. The method is implemented in Python and demonstrated on a large-scale dataset, I-24 MOTION INCEPTION. New insights revealed from this demonstration with emerging phenomena include: (a) we demonstrate that waves do generate, propagate, and dissipate at a scale (miles and hours) and ubiquity never observed before; (b) wave fronts and tails travels at a consistent speed for a critical speed between 10-20 mph, with propagation variation across lanes, where wave speed on the outer lane are less consistent compared to those on the inner lane; (c) wave fronts and tails propagate at different speeds; (d) wave boundaries capture rich and non-trivial wave topologies (with several merges and bifurcations per wave boundary on average), highlighting the complexity of waves.
@@
stop-and-go waves , traffic flow , freeway traffic , trajectory data , connected component , graph theory
@@ 1.1 Motivation and challenges
Stop-and-go waves have garnered considerable attention from the fields of mathematics, physics, and traffic engineering, and remains a persistent and complex challenge. Since their first reporting in the Lincoln Tunnel experiments [1; 2], understanding stop-and-go waves has become one of the focal points of traffic science. Empirical data allows us to observe and analyze stop-go-wave patterns; though many legacy datasets [3; 4; 5] enable these efforts, the field of traffic science continues to suffer from a paucity of comprehensive datasets [6] encompassing a broader spectrum of traffic wave phenomena, which are essential for deriving more profound insights through empirical analysis.
In recent years, advancements in computer vision technologies have facilitated the expansion of large-scale traffic trajectory datasets, such as pNEUMA drone data [7], Zen Traffic Data [8] and I-24 MOTION data [9], which can provide million or even billions of trajectory points. The field of trajectory data analysis has entered a new era of massive data [7] and brings up new research opportunities [6]. Scaling up the analysis to handle massive datasets has been both a clear research need and a significant challenge. In the context of stop-and-go wave analysis, it is feasible to manually label each wave in the NGSIM dataset and handle each case individually, as only a limited number of waves are observed. However, when the number of waves scales up to hundreds or even thousands, the necessity for an automated method to segment wave components becomes critical — a task that, to the best of our knowledge, has not been accomplished before. This motivates us to develop a method and release a tool designed to analyze large-scale datasets with massive trajectory data, thereby accelerating the traffic flow research and lowering the barrier for conducting the analyses.
@@
Moreover, large-scale data collection also reveals a broader range of complex and rich phenomena associated with these waves. The oscillatory dynamics of stop-and-go waves are intricately complex, involving stages of precursor, growth, stable, and decay [10]. Empirically, the wave propagation speed is consistent [11] and displays a concave growth pattern [12]. Wave patterns, including characteristics such as wave speed and wave duration, are documented and summarized in the literature [3; 4; 13]. It has been observed that waves can split [14], leading to bifurcation and merging [15] behaviors within wave dynamics. However, these phenomena have not been thoroughly explored, primarily due to the limited observations from empirical data, and existing methods are not designed to effectively capture or analyze bifurcation and merging, particularly in terms of the wave topology. A more detailed review of the characteristics observed in both experimental and empirical data is provided in Section 2. As shown in Figure 1a, multiple waves propagate diagonally, showcasing the complexity and richness of dynamic behaviors such as wave bifurcation and merging, which form phenomena resembling a “river delta” in the space-time diagram. This growing complexity underscores the need for scalable methods to capture the topology of waves, which could significantly accelerate research into the nature of traffic waves. To date, no existing approaches are capable of unifying all observed phenomena to provide a comprehensive understanding of the dynamics of stop-and-go waves, which motivates us to develop a method capable of analyzing these complex phenomena.
@@ 1.2 Research questions and contributions
    we propose a method for representing stop-and-go waves as a graph, incorporating the concept of connected components to describe the complex phenomena, including wave generation, propagation, dissipation, with a particular emphasis on bifurcation and merging. This method is capable of identifying both the wave fronts and tails, as well as the wave topology in stop-and-go traffic;
@@
    we implement the proposed methods in Python to support reproducible research, with data and tools will be provided at https://i24motion.org;
@@
    we apply our method to a large-scale dataset, I-24 MOTION INCEPTION, which is the first demonstration of the full life-cycle of stop-and-go waves evolving on a freeway, offering new insights into the wave dynamics with extensive data:
    1. (a)
      we demonstrate that waves do generate, propagate, and dissipate at a scale (miles and hours) and ubiquity never observed before;
    2. (b)
      wave fronts and tails travels at a consistent speed for a critical speed between 10-20 mph, with propagation variation across lanes, where wave speed on the outer lanes are less consistent compared to those on the inner lanes, which may be explained by boundary effects of on-ramps and off-ramps;
    3. (c)
      wave fronts and tails propagate at different speeds;
    4. (d)
      wave boundaries capture rich and non-trivial wave topologies (with several merges and bifurcations per wave boundary on average), highlighting the complexity of waves.
This paper is organized as follows. Section 2 presents a literature review on the stop-and-go wave analysis characteristics and its identification techniques. Section 4 outlines the method for modeling stop-and-go waves as a graph, including the use of connected components to represent these complex phenomena. Section 5 presents the data utilized in this study, along with the parameters and hyperparameters employed. Section 6 showcases the identified results, while Section 7 discusses the insights gained from the analysis. Section 8 summarize the findings and future work.
@@ 2.1 Stop-and-go wave and its characteristics
Stop-and-go waves, also called traffic oscillations [16; 17; 18], or wide-moving jammed waves [19; 20; 21; 22; 23; 24; 25; 26] or shock waves [27; 28; 29], have different terminologies in different research communities [30]. In this paper, stop-and-go waves are referred to the repeated cycles of deceleration and acceleration engaged by vehicles [13; 31; 32]. At the very beginning, inductive loop detectors were the main data sources used by researchers to analyze stop-and-go traffic [33; 34; 35; 36]. Later, the well-known NGSIM dataset [5] provided access to trajectory data, enabling the community to investigate wave phenomena on both microscopic and macroscopic scales. This made a significant contribution to the empirical analysis of stop-and-go waves, which was followed by numerous other trajectory data collection efforts worldwide [37; 8; 38; 9].
Another significant push in understanding stop-and-go waves comes from field experiments, which enable researchers to more effectively discern the causes of wave generation and propagation. [39] conducted a single-lane ring-road experiment in Japan, revealing for the first time that stop-and-go waves can occur due to instabilities in driving behavior, even in the absence of any bottlenecks. [40] later replicated the experiment under the same conditions and used one automated vehicle to dissipate the traffic waves. A research team in China [41; 15; 42; 43] also conducted multiple comprehensive field experiments with multiple-vehicle platoon on open road and closed track in different locations, showing wave features in lightly congested traffic [42], hyper-congested traffic [15] and the growing pattern of traffic oscillations [43]. In recent years, there have been numerous studies on field testing automated cruise control vehicles [44; 45; 46; 47; 48; 49; 50; 51] to gather further insights into how these vehicles impact and control stop-and-go waves.
@@
Treiterer’73 [57; 4] 209 vehicles Open freeway United States Camera
@@
From both the empirical and the experimental data, the stop-and-go waves typically have the life-cycles including generation, propagation, merge and bifurcation, and dissipation [22; 10; 61]. Below is a summary of the patterns reported in the literature:
@@
    Generation. Stop-and-go waves can be generated from various causes, including endogenous reasons like instability [39; 13], driving behaviour [14; 62] and self-organized criticality [63] and exogenous reasons including fixed bottleneck, moving bottleneck [64; 65; 66], lane-changing maneuvers [67; 68] and other disturbances or perturbations to the system [69]. Meta-stability [70; 52] is also discussed in relation to wave generation, particularly in terms of how a temporary homogeneous flow with high velocity can occur just before the waves are generated. However, the collection of empirical data is still insufficient to statistically reason through the generation of waves.
@@
    Propagation. Wave propagation is the most frequently discussed pattern and is the most intuitive to experience during daily driving [71]. One characteristic of these waves is their propagation against the traffic flow at a relatively constant speed of 10-20 km/hr, with variations depending on location [57; 72; 11; 73]. The upstream and downstream fronts of congested traffic are discussed separately in [22; 17; 18], and both warrant further investigation. At the microscopic level, the wave speed may vary over time depending on driver behavior [14]. Another characteristic observed in experimental and empirical data is that the growth of wave propagation exhibits a concave pattern, as measured by the standard deviation of individual vehicle speeds [12; 54; 43; 74]. Other measurements for wave propagation stage include duration and period [18; 32].
@@
    Merge and bifurcation. Only a few pieces of literature discuss these phenomena in detail. [15] reported that the structure of hyper-congested traffic may involve small jams merging into larger ones slowly, with larger jams occasionally breaking into smaller ones. The studies by [75; 61] considered the effect of wave-absorbing behavior in traffic control, which may lead to additional wave phenomena like wave bifurcation, this has also been observed in simulations [76; 61]. These patterns are often obscured within the internal dynamics of congested traffic, where empirical aggregated traffic measurements tend to average out these phenomena.
@@
    Dissipation. Dissipation is a frequently discussed topic in the field of control systems, examining whether stop-and-go waves can be alleviated using advanced technologies such as Connected and Automated Vehicles (CAVs) [40; 77; 78] or Intelligent Traffic Systems (ITS) [79; 80; 81]. These studies investigate the potential of these technologies to smooth traffic flow and minimize the incidence of such waves. However, from an empirical perspective, the pattern and dynamics surrounding the dissipation of these waves remain insufficiently explored. Analyses often overlook aspects such as the opposite side of the concave pattern observed during the growth phase [12], due to a lack of observations.
@@ 2.2 Stop-and-go waves identification techniques
As summarized in Section 2.1, wave generation and propagation are the most frequently discussed patterns in the literature. Consequently, most identification methods are predicated on these phenomena. Prior to the availability of trajectory data, the majority of these methods were developed using fixed sensor data as input [3; 82; 73; 17; 83], as summarized in Table 2. These models often assume that waves propagate at a constant speed, which constrains the understanding of the dynamics of stop-and-go waves. Since the release of NGSIM data, there has been a surge in the development of methods to analyze vehicle trajectory data. These methods [27; 14; 18; 31; 84; 32] have expanded the capabilities of researchers to study traffic dynamics in greater detail. In addition to detailed trajectory data, recent research has also explored the use of sparse connected vehicle trajectories to identify stop-and-go waves [85; 86]. This approach leverages limited but strategically connected data points from vehicles, offering a promising method for understanding and managing traffic dynamics with less comprehensive datasets.
@@
Clustering [27; 29] No No P
Wavelet Transform [18; 31] No No G, P
Trajectory decomposition [84; 32] No No G, P
@@
On the identification method, the central challenge lies in identifying the areas of interest across both spatial and temporal dimensions. Table 2 summarizes the method used to identify the stop-and-go waves. Using speed contour lines [3; 57] is the most straightforward approach for analyzing stop-and-go waves; however, it is inadequate for distinguishing the individual structures of each wave. Time-series data analysis [73; 88] within the time domain for each spatial or vehicle observation is another method to measure stop-and-go waves. Another mainstream method is to transform the data to the frequency domain, which allows for the separation of low-frequency and high-frequency signals [17; 18; 31], where the Wavelet Transform is viewed as the best of practice. However, all of the aforementioned methods primarily focus on wave generation and propagation, often neglecting the phenomena of wave merging and bifurcation. [14] provided a detailed behavioral analysis using asymmetric traffic theory, illustrating the wave propagation path across both space and time. This analysis revealed that the path of acceleration curves may split, potentially generating a secondary wave. This highlights the need for methods that not only detect and track waves but also stitch them together over time to support the full life-cycle wave analysis.
@@ 2.3 Emerging large-scale trajectory data
As stated by [6], more trajectory data are needed. In recent years, more trajectory data collection efforts comes from drones [89; 7; 37; 90; 91; 38], helicopter [38], and roadside infrastructures [8; 92; 59; 9]. As more data is generated, the need for tools to analyze this data and enhance research reproducibility within the community becomes more evident [93]. Additionally, with the continuous collection of data, the size of the datasets has become significantly larger than previous ones, adding more challenges for researchers to address [94]. Moreover, these datasets may uncover previously unobserved scenarios, such as the effects of automated control vehicles [38], traffic smoothing strategies [77; 95], electric vehicles [51], and traffic incidents [96]. While generating more data is a crucial first step, the development of more analytical tools is equally paramount. The majority of existing tools do not open-source their analysis software (shown in Table 2), which potentially limits subsequent analyses on different datasets. This paper aims to fill this gap by exploring the analysis of massive trajectory data to make stop-and-go wave analysis at scale possible.

### v2

@@ Abstract
Analyzing stop-and-go waves at the scale of miles and hours of data is an emerging challenge in traffic research. The past 5 years have seen an explosion in the availability of large-scale traffic data containing traffic waves and complex congestion patterns, making existing approaches unsuitable for repeatable and scalable analysis of traffic waves in these data. This paper makes a first step towards addressing this challenge by introducing an automatic and scalable stop-and-go wave identification method capable of capturing wave generation, propagation, dissipation, as well as bifurcation and merging, which have previously been observed only very rarely. Using a concise and simple critical-speed based definition of a stop-and-go wave, the proposed method identifies all wave boundaries that encompass spatio-temporal points where vehicle speed is below a chosen critical speed. The method is built upon a graph representation of the spatio-temporal points associated with stop-and-go waves, specifically wave front (start) points and wave tail (end) points, and approaches the solution as a graph component identification problem. It enables the measurement of wave properties at scale. The method is implemented in Python and demonstrated on a large-scale dataset, I-24 MOTION INCEPTION. Our results show insights on the complexity of traffic waves. Traffic waves can bifurcate and merge at a scale that has never been observed or described before. The clustering analysis of all the identified wave components reveals the different topological structures of traffic waves. We explored that the wave merge or bifurcation points can be explained by spatial features. The gallery of all the identified wave topologies is demonstrated at https://trafficwaves.github.io/.
@@
stop-and-go wave , traffic flow , freeway traffic , trajectory data , graph theory , connected component
@@ 1.1 Motivation and challenges
Stop-and-go waves have garnered considerable attention from the fields of mathematics, physics, and traffic engineering, and remains a persistent and complex challenge. Since their first reporting in the Holland and Lincoln Tunnel experiments [1; 2; 3; 4], understanding stop-and-go waves has become one of the focal points of traffic science. Empirical data allows us to observe and analyze stop-go-wave patterns; though many legacy datasets [4; 5; 6] enable these efforts, the field of traffic science continues to suffer from a paucity of comprehensive datasets [7] encompassing a broader spectrum of traffic wave phenomena, which are essential for deriving more profound insights through empirical analysis.
In recent years, advancements in computer vision technologies have facilitated the expansion of large-scale traffic trajectory datasets, such as pNEUMA drone data [8], Zen Traffic Data [9], DLR highway traffic (DLR-HT) data [10] and I-24 MOTION data [11], which can provide million or even billions of trajectory points. The field of trajectory data analysis has entered a new era of massive data [8] and brings up new research opportunities [7]. Scaling up the analysis to handle massive datasets has been both a clear research need and a significant challenge. In the context of stop-and-go wave analysis, it is feasible to manually label each wave in the NGSIM dataset and handle each case individually, as only a limited number of waves are observed. However, when the number of waves scales up to hundreds or even thousands, the necessity for an automated method to segment wave components becomes critical — a task that, to the best of our knowledge, has not been accomplished before. This motivates us to develop a method and release a tool designed to analyze large-scale datasets with massive trajectory data, thereby accelerating the traffic flow research and lowering the barrier for conducting the analyses.
@@
Moreover, large-scale data collection also reveals a broader range of complex and rich phenomena associated with these waves. The oscillatory dynamics of stop-and-go waves are intricately complex, involving stages of precursor, growth, stable, and decay [12]. Empirically, the wave propagation speed is consistent [13] and displays a concave growth pattern [14]. Wave patterns, including characteristics such as wave speed and wave duration, are documented and summarized in the literature [4; 5; 15]. It has been observed that waves can split [16], leading to bifurcation and merging [17] behaviors within wave dynamics. However, these phenomena have not been thoroughly explored, primarily due to the limited observations from empirical data, and existing methods are not designed to effectively capture or analyze bifurcation and merging, particularly in terms of the wave topology. A more detailed review of the characteristics observed in both experimental and empirical data is provided in Section 2. As shown in Figure 1a, multiple waves propagate diagonally, showcasing the complexity and richness of dynamic behaviors such as wave bifurcation and merging, which form phenomena resembling a “river delta” in the space-time diagram. This growing complexity underscores the need for scalable methods to capture the topology of waves, which could significantly accelerate research into the nature of traffic waves. To date, no existing approaches are capable of unifying all observed phenomena to provide a comprehensive understanding of the dynamics of stop-and-go waves, which motivates us to develop a method capable of analyzing these complex phenomena.
@@ 1.2 Research questions and contributions
    We propose a graph representation of stop-and-go waves, capable of capturing the complex traffic wave topological phenomena, including wave generation, propagation, dissipation, bifurcation, and merging.
@@
    We develop an automatic approach for identifying and measuring stop-and-go wave boundaries and topologies at scale with implementation to a large-scale dataset.
@@
    Our results show insights on the complexity of traffic waves. Traffic waves can bifurcate and merge at a scale that has never been observed or described before. The clustering analysis of all the identified wave components reveals the different topological structures of traffic waves. We explored that the wave merge or bifurcation points can be explained by spatial features. The gallery of all the identified wave topologies is demonstrated at https://trafficwaves.github.io/.
This paper is organized as follows. Section 2 presents a literature review on the stop-and-go wave analysis characteristics and its identification techniques. Section 4 outlines the method for modeling stop-and-go waves as a graph, including the use of connected components to represent these complex phenomena. Section 5 presents the data utilized in this study, along with the parameters and hyperparameters employed. Section 6 showcases the identified results. Section 7 discusses the statistics of traffic wave measurements and explore the insights that graph topological analysis offers to traffic modeling. Section 8 summarize the findings and future work.
@@ 2.1 Stop-and-go wave and its characteristics
Stop-and-go waves, also called traffic oscillations [18; 19; 20], or wide-moving jammed waves [21; 22; 23; 24; 25; 26; 27; 28] or shock waves [29; 30; 31], have different terminologies in different research communities [32]. In this paper, stop-and-go waves are referred to the repeated cycles of deceleration and acceleration engaged by vehicles [15; 33; 34]. At the very beginning, inductive loop detectors were the main data sources used by researchers to analyze stop-and-go traffic [35; 36; 37; 38]. Later, the well-known NGSIM dataset [6] provided access to trajectory data, enabling the community to investigate wave phenomena on both microscopic and macroscopic scales. This made a significant contribution to the empirical analysis of stop-and-go waves, which was followed by numerous other trajectory data collection efforts worldwide [39; 9; 40; 11].
Another significant push in understanding stop-and-go waves comes from field experiments, which enable researchers to more effectively discern the causes of wave generation and propagation. [41] conducted a single-lane ring-road experiment in Japan, revealing for the first time that stop-and-go waves can occur due to instabilities in driving behavior, even in the absence of any bottlenecks. [42] later replicated the experiment under the same conditions and used one automated vehicle to dissipate the traffic waves. A research team in China [43; 17; 44; 45] also conducted multiple comprehensive field experiments with multiple-vehicle platoon on open road and closed track in different locations, showing wave features in lightly congested traffic [44], hyper-congested traffic [17] and the growing pattern of traffic oscillations [45]. In recent years, there have been numerous studies on field testing automated cruise control vehicles [46; 47; 48; 49; 50; 51; 52; 53] to gather further insights into how these vehicles impact and control stop-and-go waves.
@@
Treiterer’73 [59; 5] 209 vehicles Open freeway United States Camera
@@
MiTra [63] 2.1hr $\times$ 0.9km Open freeway Germany Camera
@@
From both the empirical and the experimental data, the stop-and-go waves typically have the life-cycles including generation, propagation, merge and bifurcation, and dissipation [24; 12; 64; 65]. Below is a summary of the patterns reported in the literature:
@@
    Generation. Stop-and-go waves can be generated from various causes, including endogenous reasons like instability [41; 15], driving behaviour [16; 66] and self-organized criticality [67] and exogenous reasons including fixed bottleneck, moving bottleneck [68; 69; 70], lane-changing maneuvers [71; 72] and other disturbances or perturbations to the system [73]. Meta-stability [74; 54] is also discussed in relation to wave generation, particularly in terms of how a temporary homogeneous flow with high velocity can occur just before the waves are generated. However, the collection of empirical data is still insufficient to statistically reason through the generation of waves.
@@
    Propagation. Wave propagation is the most frequently discussed pattern and is the most intuitive to experience during daily driving [75]. One characteristic of these waves is their propagation against the traffic flow at a relatively constant speed of 10-20 km/hr, with variations depending on location [59; 76; 13; 77]. The upstream and downstream fronts of congested traffic are discussed separately in [24; 19; 20], and both warrant further investigation. At the microscopic level, the wave speed may vary over time depending on driver behavior [16]. Another characteristic observed in experimental and empirical data is that the growth of wave propagation exhibits a concave pattern, as measured by the standard deviation of individual vehicle speeds [14; 56; 45; 78]. Other measurements for wave propagation stage include duration and period [20; 34].
@@
    Merge and bifurcation. Only a few pieces of literature discuss these phenomena in detail. [17] reported that the structure of hyper-congested traffic may involve small jams merging into larger ones slowly, with larger jams occasionally breaking into smaller ones. The studies by [79; 64] considered the effect of wave-absorbing behavior in traffic control, which may lead to additional wave phenomena like wave bifurcation, this has also been observed in simulations [80; 64]. These patterns are often obscured within the internal dynamics of congested traffic, where empirical aggregated traffic measurements tend to average out these phenomena.
@@
    Dissipation. Dissipation is a frequently discussed topic in the field of control systems, examining whether stop-and-go waves can be alleviated using advanced technologies such as Connected and Automated Vehicles (CAVs) [42; 81; 82] or Intelligent Traffic Systems (ITS) [83; 84; 85]. These studies investigate the potential of these technologies to smooth traffic flow and minimize the incidence of such waves. However, from an empirical perspective, the pattern and dynamics surrounding the dissipation of these waves remain insufficiently explored. Analyses often overlook aspects such as the opposite side of the concave pattern observed during the growth phase [14], due to a lack of observations.
@@ 2.2 Stop-and-go waves identification techniques
As summarized in Section 2.1, wave generation and propagation are the most frequently discussed patterns in the literature. Consequently, most identification methods are predicated on these phenomena. Prior to the availability of trajectory data, the majority of these methods were developed using fixed sensor data as input [4; 86; 77; 19; 87], as summarized in Table 2. These models often assume that waves propagate at a constant speed, which constrains the understanding of the dynamics of stop-and-go waves. Since the release of NGSIM data, there has been a surge in the development of methods to analyze vehicle trajectory data. These methods [29; 16; 20; 33; 88; 34; 89] have expanded the capabilities of researchers to study traffic dynamics in greater detail. In addition to detailed trajectory data, recent research has also explored the use of sparse connected vehicle trajectories to identify stop-and-go waves [90; 91; 89]. This approach leverages limited but strategically connected data points from vehicles, offering a promising method for understanding and managing traffic dynamics with less comprehensive datasets.
@@
Clustering [29; 31; 93] No No P
Wavelet Transform [20; 33] No No G, P
Trajectory decomposition [88; 34] No No G, P
@@
On the identification method, the central challenge lies in identifying the areas of interest across both spatial and temporal dimensions. Table 2 summarizes the method used to identify the stop-and-go waves. Using speed contour lines [4; 59] is the most straightforward approach for analyzing stop-and-go waves; however, it is inadequate for distinguishing the individual structures of each wave. Time-series data analysis [77; 94] within the time domain for each spatial or vehicle observation is another method to measure stop-and-go waves. Another mainstream method is to transform the data to the frequency domain, which allows for the separation of low-frequency and high-frequency signals [19; 20; 33], where the Wavelet Transform is viewed as the best of practice. However, all of the aforementioned methods primarily focus on wave generation and propagation, often neglecting the phenomena of wave merging and bifurcation. [16] provided a detailed behavioral analysis using asymmetric traffic theory, illustrating the wave propagation path across both space and time. This analysis revealed that the path of acceleration curves may split, potentially generating a secondary wave. This highlights the need for methods that not only detect and track waves but also stitch them together over time to support the full life-cycle wave analysis.
@@ 2.3 Emerging large-scale trajectory data
As stated by [7], more trajectory data are needed. In recent years, more trajectory data collection efforts comes from drones [95; 8; 39; 96; 97; 40], helicopter [40], and roadside infrastructures [9; 98; 61; 11]. As more data is generated, the need for tools to analyze this data and enhance research reproducibility within the community becomes more evident [99]. Additionally, with the continuous collection of data, the size of the datasets has become significantly larger than previous ones, adding more challenges for researchers to address [100]. Moreover, these datasets may uncover previously unobserved scenarios, such as the effects of automated control vehicles [40], traffic smoothing strategies [81; 101], electric vehicles [53], and traffic incidents [102]. While generating more data is a crucial first step, the development of more analytical tools is equally paramount. The majority of existing tools do not open-source their analysis software (shown in Table 2), which potentially limits subsequent analyses on different datasets. This paper aims to fill this gap by exploring the analysis of massive trajectory data to make stop-and-go wave analysis at scale possible.